For equilibrium at 25 oC with the atmosphere, what are the aqueous concentrations of N2 (mg/L), O2 (mg/L) and CO2 (mol/L) in natural water? Example: N2 : 0.79 atm O2 : 0.21 atm CO2 : 10-3.38 atm
Answer #1

Answer:-

Given:-

temperature (T) = 25 0C = 25 + 273 = 298 K

partial pressure of N2 gas = 0.79 atm

partial pressure of O2 gas = 0.21 atm

partial pressure of CO2 gas = 10-3.38 atm

concentration of N2 (mg/L) = ?

concentration of O2 (mg/L) = ?

concentration of CO2 (mol/L) = ?

As we know that

gas constant (R) = 0.08206 L.atm.K-1.mol-1

So according to the ideal gas equation

PV = nRT

or

P = nRT / V

or

P = CRT (since C = n /V )

where

C = concentration of gas in mol / L

therefore

partial pressure of CO2 gas (PCO2) = CCO2RT

concentration of CO2 (CCO2) = partial pressure of CO2 gas (PCO2) / RT

concentration of CO2 (CCO2) = 10-3.38 atm / 0.08206 L.atm.K-1.mol-1\times 298 K

concentration of CO2 (CCO2) = 10-3.38 atm / 24.45388‬ L.atm..mol-1

concentration of CO2 (CCO2) = 0.041 \times 10-3.38 mol / L (i.e the answer)

similarly

concentration of N2 (CN2) = partial pressure of N2 gas (PN2) / RT

concentration of N2 (CN2) =  0.79 atm / 0.08206 L.atm.K-1.mol-1\times 298 K

concentration of N2 (CN2) =  0.79 atm / 24.45388‬ L.atm..mol-1

concentration of N2 (CN2) = 0.032 mol / L

above concentration of N2 (CN2) shows that 0.032 mol N2 gas present in 1 L natural water.

Since we know that

1 mol of N2 = 28 g N2 = 28000 mg N2

So

1 mol of N2 = 28000 mg N2

then

0.032 mol of N2 = 0.032 \times 28000 mg N2​​​​​​​

0.032 mol of N2 = 896.0 mg N2​​​​​​​

therefore

concentration of N2 (CN2) = 0.032 mol / L = 896.0 mg / L (i.e the answer)

similarly

concentration of O2 (CO2) = partial pressure of O2 gas (PO2) / RT

concentration of O2 (CO2) =  0.21 atm / 0.08206 L.atm.K-1.mol-1\times 298 K

concentration of O2 (CO2) =  0.21 atm / 24.45388‬ L.atm..mol-1

concentration of O2 (CO2) = 0.009 mol / L

above concentration of O2 (CO2) shows that 0.009 mol O2 gas present in 1 L natural water.

Since we know that

1 mol of O2 = 32 g O2 = 32000 mg O2

So

1 mol of O2 = 32000 mg O2

then

0.009 mol of O2 = 0.009 \times 32000 mg O2

0.009 mol of O2 =   288.0 mg O2

therefore

concentration of O2 (CO2) = 0.009 mol / L = 288.0 mg / L (i.e the answer)